# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ C.7.1 PMF Representation
$=$ ${\mathbb{E}}_{G\sim\eta_{{\bm{w}}_{t-1}}(s_{t}^{\prime})}\left[{\left({1-\left|{\frac{r_{t}+\gamma G-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ $\sum_{j=0}^{K}p_{j}(s_{t}^{\prime};{\bm{w}}_{t-1})g_{j,k}(r_{t})$
$=$ $g_{K,k}(r_{t})+\sum_{j=0}^{K-1}p_{j}(s_{t}^{\prime};{\bm{w}}_{t-1})\left({g_{j,k}(r_{t})-g_{K,k}(r_{t})}\right),$
which has the same form as Eqn. (17). Following the proof of Proposition C.3 in Appendix C.4, one can show that
${\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})=\tilde{{\bm{G}}}(r_{t}){\bm{W}}^{\intercal}_{t-1}{\bm{\phi}}(s_{t}^{\prime})+\frac{1}{K+1}\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}).$ (22)
@@
${\bm{W}}_{t}=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$ (23)
$=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{W}}_{t-1}\tilde{{\bm{G}}}^{\intercal}(r_{t})-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
Compared to Linear-CTD (Eqn. (9)), this form has an additional ${\bm{C}}^{\intercal}{\bm{C}}$, and the step size is $2\iota_{K}\alpha$.
C.7.2 CDF Representation
@@
${\bm{\Theta}}_{t}={\bm{\Theta}}_{t-1}-\alpha\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right),$
where $\nabla_{{\bm{\Theta}}}$ stands for taking gradient w.r.t. ${\bm{\Theta}}_{t-1}={\bm{W}}_{t-1}{\bm{C}}^{\intercal}\in{\mathbb{R}}^{d\times K}$ in the first term ${\eta}_{{\bm{w}}_{t-1}}(s_{t})$ (the second term is regarding as a constant). One can check that $\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ is an unbiased estimate of $\nabla_{{\bm{\Theta}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{t-1}},\bm{\mathcal{T}}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right)$.
@@
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{\Theta}}_{t-1}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}}\right]$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{\Theta}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{\Theta}}_{t-1}{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r_{t}){\bm{C}}^{\intercal}-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}{\bm{C}}^{\intercal}}\right].$
@@
${\bm{\Theta}}_{t}=$ ${\bm{\Theta}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}$ (24)
$=$ ${\bm{\Theta}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{\Theta}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{\Theta}}_{t-1}{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r_{t}){\bm{C}}^{\intercal}-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}{\bm{C}}^{\intercal}}\right],$
which is equivalent to
${\bm{W}}_{t}=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}$
$=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t-1}-{\bm{\phi}}(s_{t}^{\prime})^{\intercal}{\bm{W}}_{t-1}\tilde{{\bm{G}}}^{\intercal}(r_{t})-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\intercal}}\right],$
@@
Appendix D Omitted Results and Proofs in Section 4
D.1 Proof of Lemma 5.1
@@ Appendix D Omitted Results and Proofs in Section 4
By Lemma E.1 and Eqn. (19), we have
$\left({{\mathcal{L}}({\bm{w}})}\right)^{2}=$ $W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}},\bm{\eta}_{{\bm{w}}^{\star}}}\right)$
$\leq$ $\frac{1}{1-\gamma}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}},\bm{\eta}_{{\bm{w}}^{\star}}}\right)$
$=$ $\frac{\iota_{K}}{1-\gamma}\operatorname{tr}\left({\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)}\right)$
$=$ $\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}_{1}-{\bm{\theta}}_{2}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},$

### v2

@@ D.7.1 Probability Mass Function Representation
$=$ ${\mathbb{E}}_{G\sim\eta_{{\bm{w}}_{t-1}}(s_{t+1})}\left[{\left({1-\left|{\frac{r_{t}+\gamma G-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ $\sum_{j=0}^{K}p_{j}(s_{t+1};{\bm{w}}_{t-1})g_{j,k}(r_{t})$
$=$ $g_{K,k}(r_{t})+\sum_{j=0}^{K-1}p_{j}(s_{t+1};{\bm{w}}_{t-1})\left({g_{j,k}(r_{t})-g_{K,k}(r_{t})}\right),$
which has the same form as Eqn. (20). Following the proof of Proposition D.3 in Appendix D.4, one can show that
${\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})=\tilde{{\bm{G}}}(r_{t}){\bm{W}}^{\top}_{t-1}{\bm{\phi}}(s_{t+1})+\frac{1}{K+1}\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}).$ (25)
@@
${\bm{W}}_{t}\leftarrow$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\top}{\bm{C}}^{\top}{\bm{C}}$ (26)
$=$ ${\bm{W}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{W}}_{t-1}-{\bm{\phi}}(s_{t+1})^{\top}{\bm{W}}_{t-1}\tilde{{\bm{G}}}^{\top}(r_{t})-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\top}}\right]{\bm{C}}^{\top}{\bm{C}}.$
Note that our Linear-CTD (Eqn. (13)) is equivalent to
${\bm{W}}_{t}{\leftarrow}{\bm{W}}_{t{-}1}{-}\alpha{\bm{\phi}}(s_{t})\!\!\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{W}}_{t{-}1}{-}{\bm{\phi}}(s_{t{+}1})^{\top}{\bm{W}}_{t{-}1}\tilde{{\bm{G}}}^{\top}(r_{t})-\frac{1}{K{+}1}\!\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}){-}\bm{1}_{K}}\right)^{\top}}\right],$ (27)
in the PMF representation. Compared to Eqn. (27), the SSGD (Eqn. (26)) has an additional ${\bm{C}}^{\top}{\bm{C}}$, and the step size is $2\iota_{K}\alpha$.
D.7.2 Cumulative Distribution Function Representation
@@
${\bm{\Theta}}_{t}\leftarrow{\bm{\Theta}}_{t-1}-\alpha\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}\right](s_{t})}\right),$
where $\nabla_{{\bm{\Theta}}}$ stands for taking gradient w.r.t. ${\bm{\Theta}}_{t-1}={\bm{\theta}}_{t-1}{\bm{C}}^{\top}\in{\mathbb{R}}^{d\times K}$ in the first term ${\eta}_{{\bm{\theta}}_{t-1}}(s_{t})$ (the second term is regarding as a constant). One can check that $\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}\right](s_{t})}\right)$ is an unbiased estimate of $\nabla_{{\bm{\Theta}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{\theta}}_{t-1}},\bm{\mathcal{T}}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}\right)$.
@@
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}\right](s_{t})}\right)$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s_{t})}\right)^{\top}$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{\Theta}}_{t-1}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s)}\right)^{\top}{\bm{C}}^{\top}}\right]$
$\qquad=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{\Theta}}_{t-1}-{\bm{\phi}}(s_{t+1})^{\top}{\bm{\Theta}}_{t-1}{\bm{C}}^{-\top}\tilde{{\bm{G}}}^{\top}(r_{t}){\bm{C}}^{\top}-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right].$
@@
${\bm{\Theta}}_{t}\leftarrow$ ${\bm{\Theta}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s_{t})}\right)^{\top}$
$=$ ${\bm{\Theta}}_{t-1}-2\iota_{K}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{\Theta}}_{t-1}-{\bm{\phi}}(s_{t+1})^{\top}{\bm{\Theta}}_{t-1}{\bm{C}}^{-\top}\tilde{{\bm{G}}}^{\top}(r_{t}){\bm{C}}^{\top}-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right],$
@@
D.8 Linear-CTD is mean-preserving
We will show that our Linear-CTD is mean-preserving, which was first discovered by Lyle et al. [32, Proposition 8]. In this section, we assume the first coordinate of the feature is a constant, i.e., $\phi_{1}(s)=1/\sqrt{d}$ for any $s\in{\mathcal{S}}$. As stated before, we will always assume this to ensure ${\mathscr{P}}_{{\bm{\phi}},K}$ can be uniquely defined.
Proposition D.4.
Let ${\bm{V}}_{{\bm{\theta}}}:=(V_{\bm{\theta}}(s))_{s\in{\mathcal{S}}}$ be the value function corresponding to ${\bm{\theta}}$, i.e., $V_{\bm{\theta}}(s)={\mathbb{E}}_{G\sim\eta_{{\bm{\theta}}}(s)}[G]$, then for any initialization of the Linear-TD parameter ${\bm{\psi}}_{0}$, there exists a (not unique) corresponding Linear-CTD parameter ${\bm{\theta}}_{0}$ such that ${\bm{V}}_{{\bm{\theta}}_{0}}={\bm{V}}_{{\bm{\psi}}_{0}}$, furthermore, for any $t\geq 1$ and even number $T\geq 2$, it holds that
${\bm{V}}_{{\bm{\theta}}_{t}}={\bm{V}}_{{\bm{\psi}}_{t}},\quad{\bm{V}}_{\bar{{\bm{\theta}}}_{T}}={\bm{V}}_{\bar{{\bm{\psi}}}_{T}}.$
Proof of Proposition D.4.
Recall that the updating scheme of Linear-TD is given by
${\bm{\psi}}_{t}\leftarrow$ ${\bm{\psi}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left({V_{{\bm{\psi}}_{t-1}}(s_{t})-r_{t}-\gamma V_{{\bm{\psi}}_{t-1}}(s_{t+1})}\right)$
And the updating scheme of Linear-CTD is given by
${\bm{\Theta}}_{t}\leftarrow$ ${\bm{\Theta}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s_{t})}\right)^{\intercal}.$
Let ${\bm{V}}_{{\bm{\theta}}}:=(V_{\bm{\theta}}(s))_{s\in{\mathcal{S}}}$ be the value function corresponding to ${\bm{\theta}}$, we have
$V_{\bm{\theta}}(s)=$ $\iota_{K}\sum_{k=0}^{K-1}\left({1-F_{k}(s;{\bm{\theta}})}\right)$
$=$ $\iota_{K}\left({\bm{1}_{K}-{\bm{F}}_{{\bm{\theta}}}(s)}\right)^{\intercal}\bm{1}_{K}$
$=$ $\frac{1}{2(1-\gamma)}-\iota_{K}{\bm{\phi}}(s)^{\intercal}{\bm{\Theta}}\bm{1}_{K}.$
Hence, if we take $\psi_{0,1}=\frac{\sqrt{d}}{2(1-\gamma)}$, $\psi_{0,i}=0$ for any $i\in\left\{{2,\ldots,d}\right\}$, and ${\bm{\theta}}_{0}=\bm{0}_{dK}$, it holds that
${\bm{V}}_{{\bm{\psi}}_{0}}={\bm{V}}_{{\bm{\theta}}_{0}}=\frac{1}{2(1-\gamma)}\bm{1}_{{\mathcal{S}}}.$
We can also show that for any ${\bm{\psi}}_{0}\in{\mathbb{R}}^{d}$, there exists ${\bm{\theta}}_{0}\in{\mathbb{R}}^{d\times K}$ such that ${\bm{V}}_{{\bm{\psi}}_{0}}={\bm{V}}_{{\bm{\theta}}_{0}}$. That is we need to find ${\bm{\theta}}_{0}$ such that for any $s\in{\mathcal{S}}$,
$\iota_{K}\sum_{k=0}^{K-1}{\bm{\phi}}(s)^{\intercal}{\bm{\theta}}_{0}(k)=\frac{1}{2(1-\gamma)}-{\bm{\phi}}(s)^{\intercal}{\bm{\psi}}_{0}.$
We can take ${\bm{\theta}}_{0}$ satisfying the following equations to make the above equation hold
$\iota_{K}\sum_{k=0}^{K-1}\theta_{0}(k,1)=\frac{\sqrt{d}}{2(1-\gamma)}-\psi_{0,1},$
and
$\iota_{K}\sum_{k=0}^{K-1}{\bm{\theta}}_{0}(k)_{-1}=-{\bm{\psi}}_{0,-1}.$
Furthermore, for any $t\geq 1$, we have for any $s\in{\mathcal{S}}$,
$V_{{\bm{\theta}}_{t}}(s)=$ $\frac{1}{2(1-\gamma)}-\iota_{K}{\bm{\phi}}(s)^{\intercal}{\bm{\Theta}}_{t}\bm{1}_{K}$
$=$ $\frac{1}{2(1-\gamma)}-\iota_{K}{\bm{\phi}}(s)^{\intercal}\left({{\bm{\Theta}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s_{t})}\right)^{\intercal}}\right)\bm{1}_{K}$
$=$ $V_{{\bm{\theta}}_{t-1}}(s)+\alpha\iota_{K}\left({{\bm{\phi}}(s)^{\intercal}{\bm{\phi}}(s_{t})}\right)\left({{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})-{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s_{t})}\right)^{\intercal}\bm{1}_{K}$
$V_{{\bm{\psi}}_{t}}(s)=$ ${\bm{\phi}}(s)^{\intercal}{\bm{\psi}}_{t}$
$=$ ${\bm{\phi}}(s)^{\intercal}\left({{\bm{\psi}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left({V_{{\bm{\psi}}_{t-1}}(s_{t})-r_{t}-\gamma V_{{\bm{\psi}}_{t-1}}(s_{t+1})}\right)}\right)$
$=$ $V_{{\bm{\psi}}_{t-1}}(s)-\alpha\left({{\bm{\phi}}(s)^{\intercal}{\bm{\phi}}(s_{t})}\right)\left({V_{{\bm{\psi}}_{t-1}}(s_{t})-r_{t}-\gamma V_{{\bm{\psi}}_{t-1}}(s_{t+1})}\right).$
We need to check that, if ${\bm{V}}_{{\bm{\theta}}_{t-1}}={\bm{V}}_{{\bm{\psi}}_{t-1}}$, it holds that
$\iota_{K}\left({{\bm{F}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{\theta}}_{t-1}}}}(s_{t})-{\bm{F}}_{{\bm{\theta}}_{t-1}}(s_{t})}\right)^{\intercal}\bm{1}_{K}=V_{{\bm{\psi}}_{t-1}}(s_{t})-r_{t}-\gamma V_{{\bm{\psi}}_{t-1}}(s_{t+1}),$
which is the direct corollary of the following fact
$\text{LHS}=V_{{\bm{\psi}}_{t}}(s_{t})-{\mathbb{E}}_{X\sim{\bm{\Pi}}_{K}\left({b_{r,\gamma}}\right)_{\#}\eta_{{\bm{\theta}}_{t-1}}(s_{t+1})}[X]=V_{{\bm{\psi}}_{t}}(s_{t})-r_{t}-\gamma V_{{\bm{\theta}}_{t-1}}(s_{t+1}),$
by Lemma D.2. And we can obtain ${\bm{V}}_{\bar{\bm{\theta}}_{T}}={\bm{V}}_{\bar{\bm{\psi}}_{T}}$ by using the facts that ${\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$ is an affine space, ${\mathscr{V}}_{{\bm{\phi}}}$ is a linear space and taking expectation is a linear operator. ∎
Lemma D.2.
For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$, it holds that
${\mathbb{E}}_{X\sim\nu}[X]={\mathbb{E}}_{X\sim{\bm{\Pi}}_{K}\nu}[X].$
@@ D.8 Linear-CTD is mean-preserving
By Eqn. (8), $\bm{\Pi}_{K}\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ is uniquely identified with a vector ${\bm{p}}_{\nu}=\left({p_{k}(\nu)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$, where
$p_{k}(\nu)=\int_{\left[{0,(1-\gamma)^{-1}}\right]}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}\nu(dx).$
Hence, for any $x\in[0,(1-\gamma)^{-1}]$, we define $x_{\text{lb}}:=\max\{y\in\{x_{0},\ldots,x_{K}\}\colon x\leq y\}$, then
$\sum_{k=0}^{K}x_{k}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}=$ $x_{\text{lb}}\left({1-\frac{x-x_{\text{lb}}}{\iota_{K}}}\right)+\left({x_{\text{lb}}+\iota_{K}}\right)\left({1-\frac{x_{\text{lb}}+\iota_{K}-x}{\iota_{K}}}\right)$
$=$ $x_{\text{lb}}+\iota_{K}\left({1-\frac{x_{\text{lb}}+\iota_{K}-x}{\iota_{K}}}\right)$
$=$ $x,$
therefore,
${\mathbb{E}}_{X\sim{\bm{\Pi}}_{K}\nu}[X]=$ $\sum_{k=0}^{K}x_{k}\int_{\left[{0,(1-\gamma)^{-1}}\right]}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}\nu(dx)$
$=$ $\int_{\left[{0,(1-\gamma)^{-1}}\right]}\sum_{k=0}^{K}x_{k}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}\nu(dx)$
$=$ $\int_{\left[{0,(1-\gamma)^{-1}}\right]}x\nu(dx)$
$=$ ${\mathbb{E}}_{X\sim\nu}[X].$
Appendix E Omitted Results and Proofs in Section 4
E.1 Proof of Lemma 5.1
Proof.
By Lemma G.1 and Eqn. (22), we have
$\left({{\mathcal{L}}({\bm{\theta}})}\right)^{2}=$ $W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{\theta}}},\bm{\eta}_{{\bm{\theta}}^{\star}}}\right)$
$\leq$ $\frac{1}{1-\gamma}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{\theta}}},\bm{\eta}_{{\bm{\theta}}^{\star}}}\right)$
$=$ $\frac{\iota_{K}}{1-\gamma}\operatorname{tr}\left({\left({{\bm{\Theta}}-{\bm{\Theta}}^{\star}}\right)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}\left({{\bm{\Theta}}-{\bm{\Theta}}^{\star}}\right)}\right)$
$=$ $\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}-{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},$